A car with a mass of 1,220 kg is moving around a curve with a radius of 40.2 m at a constant speed of 41.3 m/s. What is the magnitude of the force required to produce this circular motion?
Answer #1

The magnitude of foece can be calculated as,

F = mv2/r

Where F is force

m is mass in kg

v is velocity in m/s

r is radius in m

Therefore, F = 1220×(41.3)2/40.2

F = 2080941.8/40.2

F = 51764.7214 N